OpenStudy (anonymous):

5x^2-15x/5x-8 divided by x-3/5x-8 divide and simplify to lowest terms

OpenStudy (anonymous):

start with \[\frac{5x^2-15x}{5x-8}\times \frac{5x-8}{x-3}\] then cancel as much as you can

OpenStudy (anonymous):

\[\frac{5x(x-3)(5x-8)}{(5x-8)(x-3)}=\frac{5x\cancel{x-3}\cancel{5x-8}}{\cancel {5x-8}\cancel{x-3}}=5x\]

OpenStudy (anonymous):

it ends up just 5x ?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yes